The largest angle in a triangle is 5 time the measure of the smallest angle. The middle-sized angle is 20 degree greater than twice the measure of the smallest angle. Find the measure of the three angles?

Answer by ankor@dixie-net.com(22740)

The largest angle in a triangle is 5 time the measure of the smallest angle.
The middle-sized angle is 20 degree greater than twice the measure of the smallest angle.
Find the measure of the three angles?
:
Three angles, a, b, c
:
"The largest angle in a triangle is 5 time the measure of the smallest angle."
a = 5c
" The middle-sized angle is 20 degree greater than twice the measure of the smallest angle."
b = 2c + 20
we know
a + b + c = 180
replace a and b
5c + (2c+20) + c = 180
5c + 2c + c = 180 - 20
8c = 160
c = 160/8
c = 20 degrees is the smallest angle
then
a = 5(20)
a = 100 degrees is the largest
and
b = 2(20) + 20
b = 60 degrees
:
:
Check 100 + 60 + 20 = 180
Find the measure of the three angles?